Find the QR decomposition of the following matrix: [1,2,5;-1,1,4;-1,4,-3;1,-4,7;1,2,1]
Answer #1

The given matrix is: A= | 1.000 -1.000 -1.000 1.000 | 1.000 2.000 1.000 4.000 -4.000 2.000 5.000 4.000 -3.000 7.000 1.000 k =a = 2.236 B = 0.138 k = 1 2i = 2.236 0.927 3.927 -3.927 2.073 H = / 1.000 0.000 0.000 –0.155 0.000 -0.655 0.000 0.655 10.000-0.447 0.447 Q2 = 1 0.447 -0.447 1 -0.447 -0.500 -0.000 -0.500 0.500 -0.500 0.439 0.781 0.049 0.439 -0.049 -0.236 -0.549 -0.2